OpenStudy (anonymous):

PLEASE, REALLY NEED ANSWERS TO THIS s=2πr^2+2πrh for h STEP BY STEP PLEASE

jimthompson5910 (jim_thompson5910):

\[\Large s=2\pi r^2+2\pi rh\] \[\Large s-2\pi r^2=2\pi rh\] \[\Large \frac{s-2\pi r^2}{2\pi r}=h\] \[\Large h=\frac{s-2\pi r^2}{2\pi r}\]

OpenStudy (anonymous):

Thank you so much Can you also solve this one please v=s^2+1/2sh for h step by step please thank you!!

jimthompson5910 (jim_thompson5910):

\[\Large v=s^2+\frac{1}{2}sh\] \[\Large v-s^2=\frac{1}{2}sh\] \[\Large 2(v-s^2)=sh\] \[\Large \frac{2(v-s^2)}{s}=h\] \[\Large h=\frac{2(v-s^2)}{s}\]

OpenStudy (anonymous):

Thank you so much!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

h = v/s - s - 1/2s